OpenStudy (fredoneedshelp):

Graph the line y+4=-4/3(x-1) in-point slope

OpenStudy (katiebuugcx):

Is it 4/3*(x-1) or 4/(3(x-1))?

OpenStudy (fredoneedshelp):

its 4/3 (x-1)

OpenStudy (katiebuugcx):

Multiply -4/3 into the (x-1) to get y+4=(-4/3x)+(4/3) Subtract the 4 over to get y=(-4/3x)+(4/3)+4 Simplify to get y=(-4/3x)+(16/3) and then just put that equation into a graphing calculator or google to get the graph https://www.google.com/webhp?sourceid=chrome-instant&ion=1&espv=2&ie=UTF-8#q=y%3D(-4%2F3x)%2B(16%2F3)

OpenStudy (fredoneedshelp):

let me check

OpenStudy (fredoneedshelp):

yeah it was wrong